# Large Language Models and Causal Inference in Collaboration: A Comprehensive Survey

### v1

@@ Abstract
Causal inference has shown potential in enhancing the predictive accuracy, fairness, robustness, and explainability of Natural Language Processing (NLP) models by capturing causal relationships among variables. The emergence of generative Large Language Models (LLMs) has significantly impacted various NLP domains, particularly through their advanced reasoning capabilities. This survey focuses on evaluating and improving LLMs from a causal view in the following areas: understanding and improving the LLMs’ reasoning capacity, addressing fairness and safety issues in LLMs, complementing LLMs with explanations, and handling multimodality. Meanwhile, LLMs’ strong reasoning capacities can in turn contribute to the field of causal inference by aiding causal relationship discovery and causal effect estimations. This review explores the interplay between causal inference frameworks and LLMs from both perspectives, emphasizing their collective potential to further the development of more advanced and equitable artificial intelligence systems.
@@ 1 Introduction
Recently Large Language Models (LLMs) have showcased remarkable versatility across a spectrum of critical tasks. An LLM is adept at tasks such as copywriting, enhancing original sentences with their distinct style and voice, responding to knowledge base queries, generating code, solving mathematical problems, and performing classification or generation tasks tailored to user requirements. Moreover, there has been a recent expansion into multi-modal variants, such as Large Visual Language Models (LVLMs) or Large Multi-modal Language Models, which broaden their input/output capabilities to encompass various modalities. This evolution has significantly enhanced both the potential and range of applications of these models.
In this survey, our primary focus is on Transformer-based Large Language Models (LLMs). The capability of LLMs is fundamentally rooted in their inference abilities, which dictates their proficiency in comprehending, processing, and providing solutions to various inquiries, as well as their adaptability to societally impactful domains. Consequently, extensive research efforts have been dedicated to measuring and enhancing these capabilities, ranging from assessing the reasoning abilities of LLMs to scrutinizing their decision-making processes and addressing challenges such as concept alignment across different modalities and mitigating hallucination. In addition, since LLMs are trained on extensive human knowledge with billions of parameters, they sometimes face challenges in appropriately prioritizing or downplaying what they have learned in different scenarios. This can lead to issues such as domain shift, where the model’s performance degrades on data that differ from the training set, and long-tail bias, where infrequent examples are not handled as effectively.
In many instances, language tasks require not only predicting or generating text based on patterns in the data but also understanding the underlying causal mechanisms driving these patterns. Causal inference has shown great potential in improving predictive accuracy, fairness, robustness, and explainability of Natural Language Processing (NLP) models [23]. With the advent of generative LLMs, a significant transformation has occurred across various NLP fields, attracting increased research interest in applying causal inference to address LLM-related challenges and augment their functionality. Such transformation also motivates this survey to outline causal methodologies and their implementation in LLMs, emphasizing their role in enriching our comprehension and application of language models.
@@
for tree= draw=red, align=center, anchor=west, child anchor=west, grow’=0, rounded corners, font= [Causal and Large Language Models (LLMs), rotate=90, anchor=south, text width=5.5cm [Causal for LLMs, [Reasoning Capacity [Model Understanding [[109, 79, 43, 47, 1, 88, 116], fill=yellow!20] ] [Commonsense Reasoning [[27, 98, 114, 123, 62, 17], fill=yellow!20] ] [Counterfactual Reasoning [[11, 70, 50, 60], fill=yellow!20] ] ] [Fairness and Bias [[68, 2, 82, 22, 124, 95, 65, 37], fill=yellow!20] ] [Safety [[15, 9, 122, 118, 69, 56], fill=yellow!20] ] [Explanation [[35, 10, 33, 119, 55], fill=yellow!20] ] [Multi-modality [[74, 45, 49, 85, 86, 121], fill=yellow!20] ] [Benchmark [[103, 115, 36, 107, 41, 72, 40, 85, 49, 18, 27], fill=yellow!20] ] ] [LLMs for Causal [Treatment Effect Estimation [[19, 24, 113], fill=yellow!20] ] [Causal Relationships Discovery [[44, 71, 4, 5, 61, 92, 7, 91, 8, 7], fill=yellow!20] ] ] ]
Figure 1: Structure Overview
Meanwhile, this survey also aims to explore how LLMs can help with the causal inference framework. Causal inference is formally defined as an intellectual discipline that considers the assumptions, study designs, and estimation strategies that allow researchers to draw causal conclusions based on data [75]. Causal inference has three main origins: potential outcomes, graphs, and structural equations, each serving unique purposes [110]. The potential outcomes framework [75] focuses on estimating causal effects through statistical inference and treatment comparisons. Graphical models, meanwhile, excel in mapping out causal pathways and visualizing relationships, with nodes representing variables and edges denoting directional influences. In this survey, we mainly discuss Pearl’s formulation of causal graphs [76], which formalized causal graphical models for presenting conditional independence relations among random variables using directed acyclic graphs (DAGs).
We summarize how LLMs can help causal inference in its two important components, i.e., causal relationship discovery and treatment effect estimation. Determining causal relationships among variables is a fundamental step in a causal inference framework because the estimation of the causal effect of variable $A$ on variable $B$ requires causal assumptions on causal relationships of other variables with $A$ and $B$. Traditionally, researchers rely on experts with subject matter knowledge to set the ground for such causal relationships. Causal discovery methods [108] provide an alternative by discovering causal graphs from observational data. LLMs have demonstrated abilities to determine such causal relationships based on pre-trained knowledge or given text. They can also be integrated with causal discovery methods to further increase the reliability of the outcome. Estimating treatment effects is central to causal inference but is hindered by the absence of counterfactual data in many cases. Leveraging the strong counterfactual reasoning abilities of LLMs, researchers have developed various ways to generate high-quality counterfactuals to enable treatment effect estimation.
The structure of the survey is given in Figure 1. We start with an introduction to recent advancements in Large Language Models in Section 2. Then we provide an overview of causal inference methods that are used and can be used to improve LLMs in Section 3. In the first half of the paper we go through how these methods are used in various problems in the LLM community: Section 4.1 overviews how causal methods are used to measure and improve the reasoning capacity of LLM, Section 4.2 and Section 4.3 focus on the fairness and safety issues while Section 4.4 introduces how the explainability of LLM are approached by causal inference methods. We also discuss the extension to construction and development of multi-modality large models in Section 4.5. Finally, we list existing work on evaluation and benchmarking LLMs from a causal perspective in Section 4.6. In the second half of the survey, we shift to how LLM extends the boundary of causal inference. Section 5.1 explains current assumptions, limitations and bottlenecks of causal inference. Section 5.3 and Section 5.2 states current works on improving treatment effect estimation and causal discoveries. We highlight several future directions in Section 6.
2 Background of Large Language Models
@@
Large Language Models (LLMs) have transformed the way we interact with and process language, opening up new possibilities for natural language understanding, generation, and communication [34]. These models are in a constant state of evolution, continually expanding the limits of what is achievable in language processing and artificial intelligence [73]. In this paper, we mainly focused on Transformer-based LLMs, and we provide an overview of their recent progress in this session.
The major breakthrough of Large Language Models (LLM) came in 2017 when the Transformer [93] was introduced. In a groundbreaking shift, the Transformer technology emerged, mastering the art of grasping long-term linguistic connections. This innovation wasn’t just theoretical—it allowed for simultaneous training on multiple GPUs, paving the way for the creation of significantly larger models [16]. Then came a pivotal moment in 2018: the unveiling of OpenAI’s GPT-1 [78]. This was more than just another step in natural language processing (NLP); it was a leap, thanks to its transformer-based architecture. Boasting 117 million parameters, GPT-1 wasn’t merely crunching data; it was weaving contextually coherent sentences, showcasing the transformative power of transformers in redefining NLP tasks [3]. Despite its initial limitations, GPT-1 didn’t just make a mark; it laid the groundwork for a new wave of AI exploration and sparked intense competition in the realm of LLMs.
In 2020, OpenAI released GPT-3 [20], which was able to generate highly coherent and natural-sounding text [13]. It was a big deal because it showed just how awesome these huge language models could be at a bunch of different language tasks [26]. Riding on the high from GPT-3’s success, OpenAI released the next iteration of their language model, GPT-4 [2, 67, 52] with the ability to generate even more coherent and natural-sounding text. Following GPT-4’s success, Google wasn’t far behind with their Bard [66]. Amazon jazzed up Alexa [42] with some cool AI features, Huawei jumped in with their Pangu models [111] and Alibaba proposed their QWEN models [6]. Then Meta brought out this thing called LLaMA [90], which was all about the first open-source foundation models. Compared to LLaMA, LLaMA2 [90] has made more explorations in reinforcement learning from human feedback (RLHF) and developed a chat-oriented version called LLaMA-chat, which generally outperforms existing open-source models across a range of helpfulness and safety benchmarks. A large number of researchers have extended LLaMA models by either instruction tuning [57] or continual pretraining. Alpaca [89] is the first open instruct-following model fine-tuned based on LLaMA. In addition, Vicuna [77] is another popular LLaMA variant, trained upon user-shared conversations collected from ShareGPT [31].
As diverse versions of LLMs emerge, they encounter common challenges. In this survey, we demonstrate that many of these challenges can be effectively addressed through causal methods. These methods encompass enhancing the reasoning capacity of LLMs, tackling fairness concerns and mitigating potential biases, ensuring safety, and enhancing the explainability of model outputs, and their extension to multi-modality versions.
Building upon this progress, there is now a burgeoning interest in expanding the scope of these models to encompass visual data, giving rise to the emergence of Large Visual Language Models (LVLMs). These models aim to integrate the understanding of both textual and visual information, opening up new avenues for more comprehensive AI systems capable of interpreting and generating content in multimodal formats.
One of the most prevalent approaches is inserting visual features as supplementary inputs to LLMs and aligning them with textual features. This method has been adapted in several large vision-language models (LVLMs) such as MiniGPT-4 [127], LLaVA [59], Mplug-Owl [106] and so on [55, 57, 117, 96, 106, 58, 51]. In this survey, we also demonstrate that the causal methods could help concur challenges encountered in existing LVLMs as above.
3 Brief Introduction of Causal Inference
3.1 Potential Outcome Framework
@@
One of the most influential frameworks in identifying and quantifying causal effects in observational data is the potential outcomes framework [80]. The potential outcomes approach associates causality with manipulation applied to units, and compares causal effects of different treatments via their corresponding potential outcomes. Following [80], we state basic concepts in the potential outcome framework.
Unit. A unit is the atomic research object in the treatment effect study. A unit can be a physical object, a firm, a patient, a person, or a collection of objects or persons, such as a classroom or a market, at a particular time point [80]. Under the potential outcome framework, the atomic research objects at different time points are different units.

### v2

@@ Abstract
Causal inference has demonstrated significant potential to enhance Natural Language Processing (NLP) models in areas such as predictive accuracy, fairness, robustness, and explainability by capturing causal relationships among variables. The rise of generative Large Language Models (LLMs) has greatly impacted various language processing tasks. This survey focuses on research that evaluates or improves LLMs from a causal view in the following areas: reasoning capacity, fairness and safety issues, explainability, and handling multimodality. Meanwhile, LLMs can assist in causal inference tasks, such as causal relationship discovery and causal effect estimation, by leveraging their generation ability and knowledge learned during pre-training. This review explores the interplay between causal inference frameworks and LLMs from both perspectives, emphasizing their collective potential to further the development of more advanced and robust artificial intelligence systems.
Large Language Models and Causal Inference in Collaboration: A Comprehensive Survey
Xiaoyu Liu1,*, Paiheng Xu1,*, Junda Wu2, Jiaxin Yuan1, Yifan Yang1, Yuhang Zhou1, Fuxiao Liu1, Tianrui Guan 1, Haoliang Wang3, Tong Yu3, Julian McAuley2, Wei Ai1, Furong Huang1 1University of Maryland, College Park, 2University of California San Diego, 3Adobe Research Correspondence: furongh@umd.edu. * denotes equal contribution.
@@ 1 Introduction
Recently Large Language Models (LLMs) have showcased remarkable versatility across a spectrum of critical tasks. Moreover, there has been a recent expansion into multi-modal variants, such as Large Vision Language Models (VLMs) or Multi-modal Large Language Models (MLLMs), which broaden their input/output capabilities to encompass various modalities. This evolution has significantly enhanced both the potential and range of applications of these models. In this survey, our primary focus is on Transformer-based LLMs and LVLMs.
The capability of LLMs is fundamentally rooted in their inference abilities, which dictates their proficiency in comprehending, processing, and providing solutions to various inquiries, as well as their adaptability to societally impactful domains Zhao et al. (2023d). Consequently, extensive research efforts have been dedicated to measuring and enhancing these capabilities, ranging from assessing the reasoning abilities of LLMs to scrutinizing their decision-making processes and addressing challenges such as concept alignment across different modalities and mitigating hallucination. Meanwhile, causal inference has shown great potential in improving predictive accuracy, fairness, robustness, and explainability of Natural Language Processing (NLP) models Feder et al. (2022), With LLMs revolutionizing various language processing tasks, there is a growing trend in applying causal inference to address LLM-related challenges and enhance their functionality. This survey outlines causal methodologies and their implementation in LLMs, emphasizing their role in enriching our comprehension and application of language models.
Moreover, this survey also aims to explore how LLMs can help with the causal inference framework. Causal inference is formally defined as an intellectual discipline that considers the assumptions, study designs, and estimation strategies that allow researchers to draw causal conclusions based on data Pearl (2009). Causal inference has three main origins: potential outcomes, graphs, and structural equations, each serving unique purposes Zeng and Wang (2022). In this survey, we mainly discuss Pearl’s formulation of causal graphs Pearl (1998), which formalized causal graphical models for presenting conditional independence among random variables using directed acyclic graphs (DAGs).
We summarize how LLMs can help causal inference in its two important components, i.e., causal relationship discovery and treatment effect estimation. Estimating causal effects between variables requires assumptions about their relationships with other variables, traditionally provided by experts LLMs, leveraging pre-trained knowledge, can assist in identifying these relationships and enhance causal discovery methods Zanga et al. (2022) for more reliable outcomes. Additionally, estimating treatment effects is often hindered by the absence of counterfactual data. By utilizing LLMs’ strong generative abilities, researchers have developed various ways to generate high-quality counterfactuals to enable treatment effect estimation. Figure 1 shows an overview of selected topics about the interplay between causal inference frameworks and LLMs from both perspectives.
@@
for tree= draw=red, align=center, anchor=west, child anchor=west, grow’=0, rounded corners, font= [Causal and Large Language Models (LLMs), rotate=90, anchor=south, text width=5.5cm [Causal for LLMs, [Reasoning Capacity [Model Understanding [Zečević et al. (2023); Romanou et al. (2023); Kim et al. (2023); Li and Li (2023)
Abdali et al. (2023); Tang et al. (2023); Zhang et al. (2023c); Chen et al. (2024b), fill=yellow!20] ] [Commonsense Reasoning [Gao et al. (2023); Willig et al. (2022); Zhang et al. (2022)
Zheng et al. (2023); Lu et al. (2022); Chen et al. (2023a), fill=yellow!20] ] [Counterfactual Reasoning [Betti et al. (2023); Miao et al. (2023); Li et al. (2023c); Liu et al. (2023d), fill=yellow!20] ] ] [Fairness and Bias [Meade et al. (2021); Achiam et al. (2023); Stanovsky et al. (2019); Ding et al. (2022); Zhou et al. (2023a); Wang et al. (2023)
Madhavan et al. (2023); Jenny et al. (2023); Xia et al. (2024); Wu et al. (2024a); Gallegos et al. (2024), fill=yellow!20] ] [Safety [Cao et al. (2022); Bao et al. (2021); Zhao et al. (2023e, 2022); Meng et al. (2021); Liu et al. (2023a), fill=yellow!20] ] [Explanation [Hou et al. (2023); Belrose et al. (2023); Gurnee et al. (2023); Zhao et al. (2023a); Liu et al. (2023b); Vig et al. (2020), fill=yellow!20] ] [Multi-modality [Pawlowski et al. (2023); Ko et al. (2023); Li et al. (2023b); Su et al. (2023); Tai et al. (2023); Zhao et al. (2023c); Chen et al. (2024a), fill=yellow!20] ] [Benchmark [Xie et al. (2023); Zhang et al. (2023b); Huang et al. (2023); Yu et al. (2023); Jin et al. (2023a); Nie et al. (2023)
Jin et al. (2023b); Su et al. (2023); Li et al. (2023b); Chen et al. (2023b); Gao et al. (2023); Chen et al. (2024b, a), fill=yellow!20] ] ] [LLMs for Causal [Treatment Effect Estimation [Chen et al. (2023c); Feder et al. (2023); Zhang et al. (2023a), fill=yellow!20] ] [Causal Relationships Discovery [Kıcıman et al. (2023); Naik et al. (2023); Antonucci et al. (2023); Long et al. (2023); Vashishtha et al. (2023)
Arsenyan and Shahnazaryan (2023); Ban et al. (2023a); Tu et al. (2023); Ban et al. (2023b, a)
Joshi et al. (2024); Feng et al. (2024); Li et al. (2024); Khatibi et al. (2024); Wan et al. (2024), fill=yellow!20] ] ] ]
Figure 1: An overview of the interplay between causal inference frameworks and LLMs.
2 Causal Inference and Large Language Models
In this section, we provide a brief introduction to both large language models (LLMs) and causal inference, laying the groundwork for understanding their interaction. We begin by outlining the core principles and recent advancements in LLMs, which have reshaped natural language processing by enabling sophisticated language generation and comprehension. Following this, we introduce the fundamental concepts of causal inference, emphasizing its role in discerning cause-and-effect relationships within data. By establishing a foundational understanding of these two fields, we aim to clarify how their integration can lead to significant advancements, enhancing both the reasoning capabilities of LLMs and the development of causal inference methodologies.
2.1 Evolution of the Large Language Models
@@
Large Language Models (LLMs) have transformed the way we interact with and process language, opening up new possibilities for natural language understanding, generation, and communication Zhao et al. (2023d). In this paper, we mainly focus on Transformer Vaswani et al. (2017) based LLMs, and we provide an overview of their recent progress in this section.
The introduction of the highly parallelizable Transformer architecture Vaswani et al. (2017), which leverages self-attention mechanisms, has led to the development of a range of Pre-trained Language Models (PLMs) with varying architectures and pre-training strategies, e.g., BERT Devlin et al. (2019), GPT2 Radford et al. (2019), and so on Radford et al. (2018); Liu et al. (2019); Lewis et al. (2020). These models largely improved performances for NLP tasks with a learning paradigm of general-purpose pre-training and task-specific fine-tuning. Researchers then find that scaling the size of the model or pre-training data often leads to further improved model capacity on downstream tasks Kaplan et al. (2020); Wei et al. (2022a); Schaeffer et al. (2024). These larger-scaled PLMs are referred to as large language models (LLMs) Brown et al. (2020); Chowdhery et al. (2022); Team et al. (2023); Achiam et al. (2023); Bai et al. (2023); Touvron et al. (2023); Dubey et al. (2024).
These LLMs demonstrated strong capacities for language processing and solving complex tasks through text generation. We briefly introduce three common strategies that are used to further enhance their performance: In-context Learning (ICL) Brown et al. (2020), which allows LLMs with natural language instruction and/or several task demonstrations (i.e., input-output pairs); Chain-of-Through (CoT) prompting Wei et al. (2024), which includes intermediate reasoning steps into prompts to guide the model toward the final answer; and Instruction tuning Sanh et al. (2024); Ouyang et al. (2022); Thoppilan et al. (2022), which refers to fine-tuning the LLMs with a mixture of multi-task datasets formatted as natural language instructions. Additionally, researchers have extended these LLMs to handle images as input Team et al. (2023); Achiam et al. (2023), leading to the development of Large Vision Language Models (VLMs) or Multi-modal Large Language Models (MLLMs). In this survey, we primarily focus on works that adopt a causal perspective to improve LLMs in their generative capacities.
2.2 Introduction of Causal Inference
In this section, we present the background knowledge of causal inference, including task descriptions, basic concepts and notations, and general solutions. More details can be found in Appendix A.
@@ 2.2 Introduction of Causal Inference
Causal inference aims to estimate the causal relationship among variables. The variables of interest are referred to as treatment, naturally, the effects of treatments are referred to as treatment effects. Ideally, the treatment effect can be measured as follows: applying different treatments to the same cohort, and then the difference in the effect is the treatment effect. However, it is impracticable for perfectly controlled experiments in many cases, requiring estimation of treatment effect from observational data. One of the most influential frameworks in identifying and quantifying causal effects in observational data is the potential outcomes framework Rubin (1974). The potential outcomes approach associates causality with manipulation applied to units, and compares causal effects of different treatments via their corresponding potential outcomes.